# Interpretation of a Discrete de Rham method as a Finite Element System

### v1

@@ 2 Lowest order case
$\inf_{v\in\mathrm{d}Y^{k-1}(T),w\in E}\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v\rangle+\langle u,w\rangle|}{\|u\|(\|v\|+\|w\|)}>0.$ (27)
@@
$\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v_{n}\rangle+\langle u,w_{n}\rangle|}{\|u\|}\to 0.$ (28)
@@
$\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v_{n}\rangle|}{\|u\|}\to 0.$ (29)
@@
$\sup_{v\in X_{0}^{k-1}(T)}\frac{|\int_{T}u_{n}\cdot\mathrm{d}v|}{|v|}\leq\sup_{w\in F}\frac{|\int_{T}u_{n}\cdot\mathrm{d}w|}{|w|}.$ (32)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v=0.$ (33)
@@
It was proved in [10, Preprint, Proposition 3.9] that $\wedge:\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)\times\mathcal{P}_{\!-}^{s}\Lambda^{l}(T)\to\mathcal{P}_{\!-}^{r+s}\Lambda^{k+l}(T)$. A simpler proof based on the Koszul operator was then provided in [3, Equation (3.16)]. Based on this result, if $u\in\mathcal{P}_{\!-}^{1}\Lambda^{k}(T)$ and $b^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k+1}(T)$ we have that $u\wedge\kappa_{T}b$ is a linear $d$-form. Its integral can therefore be obtained from evaluation at the barycentre of $T$. Since furthermore $\kappa_{T}$ was constructed with respect to this point, we get that $\int_{T}u\wedge\kappa_{T}b^{\prime}=0$. Similarly $\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0$ for $a^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k}(T)$. ∎
@@
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=\int_{\partial T}u\wedge\kappa_{T}a^{\prime}+\int_{T}u\wedge\frac{(-1)^{k+1}}{d-k}a^{\prime}.$ (34)
@@
$\int_{T^{\prime}}u\wedge\kappa_{T}a^{\prime}.$ (35)
@@
$\int_{T^{\prime}}u\wedge\kappa_{T^{\prime}}a^{\prime}=0.$ (36)
@@
For a given $u\in A^{k}(T)$ one can then compute the $\mathrm{L}^{2}$ projection of $u$ onto constants $\mathcal{P}^{0}\Lambda^{k}(T)$, from its degrees of freedom. We denote it as $\gamma_{T}u$. In practice one computes $\gamma_{T^{\prime}}u$ starting with $\dim T^{\prime}=k$ and increasing dimension of $T^{\prime}$ by one at each step.
Following VEM practice one introduces a symmetric bilinear form $s_{h}$ on $A^{k}(T)$, defined from the degrees of freedom, which is bounded and coercive in $\mathrm{L}^{2}$, uniformly with respect to $T$. It should scale like the $\mathrm{L}^{2}$ inner product (under scaling maps $x\to h_{T}x$, where $h_{T}$ is the diameter of $T$) but no consistency is required. One then defines the discrete $\mathrm{L}^{2}$-product on $A^{k}(T)$ by:
$\langle u,u^{\prime}\rangle_{h}=\int_{T}\gamma_{T}u\cdot\gamma_{T}u^{\prime}+s_{h}(u-\gamma_{T}u,u^{\prime}-\gamma_{T}u^{\prime}).$ (37)
The second term on the right side is referred to as the stablizing term.
The crucial points are that this discrete inner product is computable, coercive and consistent. It is consistent in the sense that if at least one of $u,u^{\prime}$ is constant, it gives the exact value:
$\langle u,u^{\prime}\rangle_{h}=\int_{T}u\cdot u^{\prime}.$ (38)
Remark also that the exterior derivative $\mathrm{d}:A^{k}(T)\to A^{k+1}(T)$ is computable from the degrees of freedom on $A^{k}(T)$ to those of $A^{k+1}(T)$. From there, various numerical methods for various PDEs can be defined and analyzed as in VEM.
@@ 3 High order case
$U^{k}(T)=\{u\in\mathcal{P}_{\!-}^{r+1}\Lambda^{k}(T)\ :\ u\perp\mathcal{P}^{r}\Lambda^{k}(T)\}.$ (39)
@@
$U^{0}(T)=\{u\in\mathcal{P}^{r+1}(T)\ :\ u\perp\mathcal{P}^{r}(T)\}.$ (40)
In particular this space does not contain any constant function. We also notice that for $d=\dim T$:
$U^{d}(T)=0.$ (41)
@@
  • The decomposition (42) is $\mathrm{L}^{2}$-orthogonal.
@@
    $\left\{\begin{array}[]{rcl}\displaystyle U^{k-1}\times U^{k}&\to&W^{k}(T),\\ \displaystyle(u^{\prime},u)&\mapsto&\mathrm{d}u^{\prime}+u.\end{array}\right$. (43)
@@
$\mathcal{P}_{\!-}^{r+1}\Lambda^{k}(T)=\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)\oplus W^{k}(T).$ (44)
@@
Referring to [3, Theorem 3.1], we write, for $u\in\mathcal{P}^{r+1}\Lambda^{k}(T)$:
@@
$(k+r+1)u=\mathrm{d}\kappa_{T}u+\kappa_{T}\mathrm{d}u-Q_{r}u.$ (46)
@@
$v=u+v^{\prime},\textrm{with }u\in U^{k}\textrm{ and }v^{\prime}\in\mathcal{P}^{r}\Lambda^{k}(T).$ (47)
@@
$\kappa_{T}v^{\prime}=u^{\prime}+v^{\prime\prime},\textrm{with }u^{\prime}\in U^{k-1}\textrm{ and }v^{\prime\prime}\in\mathcal{P}^{r}\Lambda^{k-1}(T).$ (49)
@@
$(k+r)v^{\prime}=\mathrm{d}u^{\prime}+\mathrm{d}v^{\prime\prime}+\kappa_{T}\mathrm{d}v^{\prime}-Q_{r-1}v^{\prime}.$ (50)
The first term on the right side is in $\mathrm{d}U^{k-1}$ and the three others are in $\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$. Inserting this in (47) we are done. ∎
@@
$Z^{k}_{0}(T)=\{u\in X^{k}_{0}(T)\ :\ \forall w\in W^{d-k}(T)\quad\int_{T}u\wedge w=0\}.$ (51)
@@
$G^{k}(T)=\{(f,g)\in\mathcal{P}^{r-1}\Lambda^{k}(T)\times\mathcal{P}^{r-1}\Lambda^{k-1}(T)\ :\ \mathrm{d}^{\star}f=0,\ \mathrm{d}^{\star}g=0\}.$ (53)
@@
$(\mathrm{d}^{\star}\mathrm{d}u,\mathrm{d}^{\star}u)\in G^{k}(T).$ (54)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge a+\int_{T}u^{\prime}\wedge b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (55)
$\int_{T}\mathrm{d}u\wedge a^{\prime}=0,\ \int_{T}u\wedge b^{\prime}$ $=0,$ (56)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime},$ (57)
@@
$\mathrm{d}^{\star}\mathrm{d}u+\mathrm{d}v$ $=f,$ (58)
$\mathrm{d}^{\star}u$ $=g.$ (59)
@@
There, we proved that the spaces of solutions to these PDEs define a FES which is compatible in the sense of [20, Definition 5.12] and that the spaces $\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$ give unisolvent degrees of freedom.
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge a$ $=\int_{T}f\cdot u^{\prime},$ (60)
$\int_{T}\mathrm{d}u\wedge a^{\prime}$ $=0.$ (61)
@@
$\int_{T}u^{\prime}\wedge b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (63)
$\int_{T}u\wedge b^{\prime}$ $=0,$ (64)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime},$ (65)
@@
Let $\overline{u}$ be the $\mathrm{L}^{2}$ projection of $u$ on $\mathcal{P}^{r}\Lambda^{d}(T)$. Due to the second equation (64) we have:
$\int_{T}(u-\overline{u})\wedge b^{\prime}=0,$ (66)
@@
Since $\mathrm{d}^{\star}g=0$, the last equation (65) is actually true for any $v^{\prime}\in Z^{d-1}_{0}(T)$. Using it for $v^{\prime}=t$ we get:
$\int_{T}|\mathrm{d}t|^{2}=\int_{T}g\cdot t=\pm\int_{T}t\wedge{\scriptstyle\bigstar}g=0.$ (68)
@@
For $k=d$ the system (55) can be solved iff $f=0$. In that case $u$ is determined up to a constant.
@@
$\pm{\scriptstyle\bigstar}b+\mathrm{d}v=f.$ (69)
@@
$A^{k}(T)=\{u\in X^{k}(T)\ :\ \forall T^{\prime}\subseteq T\quad u|_{T^{\prime}}\in B^{k}(T^{\prime})\}.$ (70)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}=\int_{T}f\cdot u^{\prime}.$ (71)
@@
Given $u\in\mathcal{P}^{r}\Lambda^{k}(T)$ we define $f=\mathrm{d}^{\star}\mathrm{d}u$ and $g=\mathrm{d}^{\star}u$. Then the PDE system is solved with zero Lagrange multipliers. ∎
@@
For $k<d$ the map $G^{k}(T)\to A^{k}_{0}(T)$ obtained from solving (55) with homogeneous boundary conditions for $u$ is injective.
@@
Suppose that the solution to (55) with $u|_{\partial T}=0$ is $u=0$. We then get:
$\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (72)
$0$ $=\int_{T}g\cdot v^{\prime},$ (73)
@@
$\forall w^{\prime}\in W^{d-k}(T)\quad\int_{T}u^{\prime}\wedge w^{\prime}=0,\textrm{i.e. }u^{\prime}\perp{\scriptstyle\bigstar}W^{d-k}(T).$ (76)
This implies that $f\in{\scriptstyle\bigstar}W^{d-k}(T)$. But in view of (44) we have ${\scriptstyle\bigstar}W^{d-k}(T)\cap\mathcal{P}^{r-1}\Lambda^{k}(T)=0$, so we get $f=0$.
@@
Any element $u\in A^{k}(T)$ is uniquely determined by the following degrees of freedom:
$\int_{T^{\prime}}u\wedge e,\quad e\in\mathcal{P}_{\!-}^{r}\Lambda^{d^{\prime}-k}(T^{\prime}),\ d^{\prime}=\dim T^{\prime}.$ (77)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (79)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime}.$ (80)
@@
$\int_{T}u\wedge e=0.$ (81)
Using (44) we then get, actually, that for all $e\in\mathcal{P}_{\!-}^{r+1}\Lambda^{d-k}(T)$:
$\int_{T}u\wedge e=0.$ (82)
@@
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=0,$ (84)
$\int_{T}g\cdot v^{\prime}$ $=\int_{T}g\cdot\mathrm{d}v^{\prime\prime}=0.$ (85)
Combining this with (80) we conclude that for all $v^{\prime}\in Z^{k-1}_{0}(T)$:
$\int_{T}u\cdot\mathrm{d}v^{\prime}=\int_{T}g\cdot v^{\prime}.$ (86)
@@
$\tilde{Z}^{k}_{0}(T)=\{u\in X^{k}_{0}(T)\ :\ \forall e\in\mathcal{P}_{\!-}^{r+1}\Lambda^{d-k}(T)\quad\int_{T}u\wedge e=0\}.$ (87)
We have $\tilde{Z}^{k}_{0}(T)\subseteq Z^{k}_{0}(T)$. The complex $\tilde{Z}^{\scriptscriptstyle{\bullet}}_{0}(T)$ is exact, again by the snake lemma and the exactness properties of $\mathcal{P}_{\!-}^{r+1}\Lambda^{{\scriptscriptstyle{\bullet}}}(T)$, including at index $k=d$.
Since $\mathrm{d}u=0$ and (82) holds we deduce that there is $t\in\tilde{Z}^{k-1}_{0}(T)$ such that $\mathrm{d}t=u$. This entails that for all $e\in\mathcal{P}_{\!-}^{r+1}\Lambda^{d-k+1}(T)$:
$\int_{T}t\wedge e=0.$ (88)
Then we have, setting $u=\mathrm{d}t$ and $v^{\prime}=t$ in (86):
$\int_{T}\mathrm{d}t\cdot\mathrm{d}t=\int_{T}g\cdot t=\int_{T}t\wedge{\scriptstyle\bigstar}g=0.$ (89)
@@
We have proved that the bilinear form (78) defines an injective map:
$A^{k}_{0}(T)\to\mathcal{P}_{\!-}^{r}\Lambda^{d-k}(T)^{\star}.$ (90)
@@
For $k<d$, we have from [16, Equation (2.28)]:
$\dim G^{k}(T)=\dim\mathcal{P}_{\!-}^{r}\Lambda^{d-k}(T).$ (91)
@@
$\dim A^{d}(T)=\dim\mathcal{P}^{r}\Lambda^{d}(T)=\dim\mathcal{P}_{\!-}^{r}\Lambda^{0}(T).$ (93)
It follows that the bilinear form (78) is invertible.
(v) Now it follows from Proposition 3.5 that the FES admits extensions in the sense of [21, Definition 2.3] (or [20, Definition page 75]). The unisolvence of the degrees of freedom follows then from a general induction argument given for FES in [21, Proposition 2.5] (or [20, Proposition 5.35]). ∎
@@
For elements $\in A^{k}(T)$ its moments against $\mathcal{P}^{r}\Lambda^{d-k}(T)$ can be computed inductively from the degrees of freedom.
@@
$(d-k+r)w=\mathrm{d}\kappa_{T}w+\kappa_{T}\mathrm{d}w-Q_{r-1}w.$ (95)
@@
$\int_{T}u\wedge\mathrm{d}\kappa_{T}w=\pm\int_{\partial T}u\wedge\kappa_{T}w\pm\int_{T}\mathrm{d}u\wedge\kappa_{T}w.$ (96)
@@
$\int_{T^{\prime}}u\wedge(\kappa_{T}-\kappa_{T^{\prime}})w+\int_{T^{\prime}}u\wedge(\kappa_{T^{\prime}}w-P_{r}\kappa_{T^{\prime}}w)+\int_{T^{\prime}}u\wedge P_{r}\kappa_{T^{\prime}}w.$ (97)
@@
In (97) the first and last terms are moments against $\mathcal{P}^{r}\Lambda^{d-1-k}(T)$, so amenable to an induction step. The middle term is zero, since it is a moment against $U^{d-1}(T^{\prime})$ representing a constraint.
@@
(iii) In the general case we remark that $\mathrm{d}\mathrm{d}u=0$. Applying the intermediate step to $\mathrm{d}u$, its moments against polynomials of degree $r$ can be computed. Going back to (96) and looking at the last term, we write:
$\int_{T}\mathrm{d}u\wedge\kappa_{T}w=\int_{T}\mathrm{d}u\wedge(\kappa_{T}w-P_{r}\kappa_{T}w)+\int_{T}\mathrm{d}u\wedge P_{r}\kappa_{T}w.$ (98)

### v2

@@ 2 Lowest order case
$\inf_{v\in\mathrm{d}Y^{k-1}(T),w\in E}\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v\rangle+\langle u,w\rangle|}{\|u\|(\|v\|+\|w\|)}>0.$ (2.26)
@@
$\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v_{n}\rangle+\langle u,w_{n}\rangle|}{\|u\|}\to 0.$ (2.27)
@@
$\sup_{u\in X^{k}_{0}(T)}\frac{|\langle u,v_{n}\rangle|}{\|u\|}\to 0.$ (2.28)
@@
$\sup_{v\in X_{0}^{k-1}(T)}\frac{|\int_{T}u_{n}\cdot\mathrm{d}v|}{|v|}\leq\sup_{w\in F}\frac{|\int_{T}u_{n}\cdot\mathrm{d}w|}{|w|}.$ (2.31)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v=0.$ (2.32)
@@
It was proved in [15, Preprint, Proposition 3.9] that $\wedge:\mathcal{P}_{\!\!-}^{r}\Lambda^{k}(T)\times\mathcal{P}_{\!\!-}^{s}\Lambda^{l}(T)\to\mathcal{P}_{\!\!-}^{r+s}\Lambda^{k+l}(T)$. A simpler proof based on the Koszul operator, rather than barycentric coordinates, was then provided in [4, Equation (3.16)]. Based on this result, if $u\in\mathcal{P}_{\!\!-}^{1}\Lambda^{k}(T)$ and $b^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k+1}(T)$ we have that $u\wedge\kappa_{T}b$ is a linear $d$-form. Its integral can therefore be obtained from evaluation at the barycentre of $T$. Since furthermore $\kappa_{T}$ was constructed with respect to this point, we get that $\int_{T}u\wedge\kappa_{T}b^{\prime}=0$. Similarly $\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0$ for $a^{\prime}\in\mathrm{P}^{0}\Lambda^{d-k}(T)$. ∎
@@
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=\int_{\partial T}u\wedge\kappa_{T}a^{\prime}+\int_{T}u\wedge\frac{(-1)^{k+1}}{d-k}a^{\prime}.$ (2.33)
@@
$\int_{T^{\prime}}u\wedge\kappa_{T}a^{\prime}.$ (2.34)
@@
$\int_{T^{\prime}}u\wedge\kappa_{T^{\prime}}a^{\prime}=0.$ (2.35)
@@
For a given $u\in A^{k}(T)$ one can then compute the $\mathrm{L}^{2}(T)$ projection of $u$ onto constants $\mathcal{P}^{0}\Lambda^{k}(T)$, from its degrees of freedom. We denote it as $\gamma_{T}u$. In practice one computes $\gamma_{T^{\prime}}u$ starting with $\dim T^{\prime}=k$ and increasing dimension of $T^{\prime}$ by one at each step.
In §4 we sketch how $\gamma_{T}$ is used as a crucial ingredient in the design of some numerical methods, that are amenable to an analysis based on Strang’s first lemma [32, Theorem 26.1 p. 192].
@@ 3 High order case
$U^{k}(T)=\{u\in\mathcal{P}_{\!\!-}^{r+1}\Lambda^{k}(T)\ :\ u\perp\mathcal{P}^{r}\Lambda^{k}(T)\}.$ (3.1)
@@
$U^{0}(T)=\{u\in\mathcal{P}^{r+1}(T)\ :\ u\perp\mathcal{P}^{r}(T)\}.$ (3.2)
In particular this space does not contain any non-zero constant function. We also notice that for $d=\dim T$:
$U^{d}(T)=0.$ (3.3)
@@
  • The decomposition (3.4) is $\mathrm{L}^{2}$-orthogonal.
@@
    $\left\{\begin{array}[]{rcl}\displaystyle U^{k-1}\times U^{k}&\to&W^{k}(T),\\ \displaystyle(u^{\prime},u)&\mapsto&\mathrm{d}u^{\prime}+u.\end{array}\right$. (3.7)
@@
$\mathcal{P}_{\!\!-}^{r+1}\Lambda^{k}(T)=\mathcal{P}_{\!\!-}^{r}\Lambda^{k}(T)\oplus W^{k}(T).$ (3.8)
@@
Referring to [4, Theorem 3.1], we write, for $u\in\mathcal{P}^{r+1}\Lambda^{k}(T)$:
@@
$(k+r+1)u=\mathrm{d}\kappa_{T}u+\kappa_{T}\mathrm{d}u-Q_{r}u.$ (3.10)
@@
$v=u+v^{\prime},\textrm{with }u\in U^{k}\textrm{ and }v^{\prime}\in\mathcal{P}^{r}\Lambda^{k}(T).$ (3.11)
@@
$\kappa_{T}v^{\prime}=u^{\prime}+v^{\prime\prime},\textrm{with }u^{\prime}\in U^{k-1}\textrm{ and }v^{\prime\prime}\in\mathcal{P}^{r}\Lambda^{k-1}(T).$ (3.13)
@@
$(k+r)v^{\prime}=\mathrm{d}u^{\prime}+\mathrm{d}v^{\prime\prime}+\kappa_{T}\mathrm{d}v^{\prime}-Q_{r-1}v^{\prime}.$ (3.14)
The first term on the right side is in $\mathrm{d}U^{k-1}$ and the three others are in $\mathcal{P}_{\!\!-}^{r}\Lambda^{k}(T)$. Inserting this in (3.11) we are done. ∎
@@
$Z^{k}_{0}(T)=\{u\in X^{k}_{0}(T)\ :\ \forall w\in W^{d-k}(T)\quad\int_{T}u\wedge w=0\}.$ (3.15)
@@
$G^{k}(T)=\{(f,g)\in\mathcal{P}^{r-1}\Lambda^{k}(T)\times\mathcal{P}^{r-1}\Lambda^{k-1}(T)\ :\ \mathrm{d}^{\star}f=0,\ \mathrm{d}^{\star}g=0\}.$ (3.17)
@@
$(\mathrm{d}^{\star}\mathrm{d}u,\mathrm{d}^{\star}u)\in G^{k}(T).$ (3.18)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge a+\int_{T}u^{\prime}\wedge b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (3.19)
$\int_{T}\mathrm{d}u\wedge a^{\prime}=0,\ \int_{T}u\wedge b^{\prime}$ $=0,$ (3.20)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime},$ (3.21)
@@
$\mathrm{d}^{\star}\mathrm{d}u+\mathrm{d}v$ $=f,$ (3.22)
$\mathrm{d}^{\star}u$ $=g.$ (3.23)
@@
There, we proved that the spaces of solutions to these PDEs define a FES which is compatible in the sense of [27, Definition 5.12] and that the spaces $\mathcal{P}_{\!\!-}^{r}\Lambda^{k}(T)$ give unisolvent degrees of freedom.
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge a$ $=\int_{T}f\cdot u^{\prime},$ (3.24)
$\int_{T}\mathrm{d}u\wedge a^{\prime}$ $=0.$ (3.25)
@@
$\int_{T}u^{\prime}\wedge b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (3.27)
$\int_{T}u\wedge b^{\prime}$ $=0,$ (3.28)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime},$ (3.29)
@@
Let $\overline{u}$ be the $\mathrm{L}^{2}$ projection of $u$ on $\mathcal{P}^{r}\Lambda^{d}(T)$. Due to the second equation (3.28) we have:
$\int_{T}(u-\overline{u})\wedge b^{\prime}=0,$ (3.30)
@@
Since $\mathrm{d}^{\star}g=0$, the last equation (3.29) is actually true for any $v^{\prime}\in Z^{d-1}_{0}(T)$. Using it for $v^{\prime}=t$ we get:
$\int_{T}|\mathrm{d}t|^{2}=\int_{T}g\cdot t=\pm\int_{T}t\wedge\raisebox{1.0pt}{$\scriptstyle\bigstar$}g=0.$ (3.32)
@@
For $k=d$ the system (3.19) can be solved iff $f=0$. In that case $u$ is determined up to a constant.
@@
$\pm\raisebox{1.0pt}{$\scriptstyle\bigstar$}b+\mathrm{d}v=f.$ (3.33)
@@
$A^{k}(T)=\{u\in X^{k}(T)\ :\ \forall T^{\prime}\subseteq T\quad u|_{T^{\prime}}\in B^{k}(T^{\prime})\}.$ (3.34)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}=\int_{T}f\cdot u^{\prime}.$ (3.35)
@@
Given $u\in\mathcal{P}^{r}\Lambda^{k}(T)$ we define $f=\mathrm{d}^{\star}\mathrm{d}u$ and $g=\mathrm{d}^{\star}u$. Then the PDE system is solved with zero Lagrange multipliers.
This holds also on the faces of $T$. ∎
@@
For $k<d$ the map $G^{k}(T)\to A^{k}_{0}(T)$ obtained from solving (3.19) with homogeneous boundary conditions for $u$ is injective.
@@
Suppose that the solution to (3.19) with $u|_{\partial T}=0$ is $u=0$. We then get:
$\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (3.36)
$0$ $=\int_{T}g\cdot v^{\prime},$ (3.37)
@@
$\forall w^{\prime}\in W^{d-k}(T)\quad\int_{T}u^{\prime}\wedge w^{\prime}=0,\textrm{ i.e. }\ u^{\prime}\perp\raisebox{1.0pt}{$\scriptstyle\bigstar$}W^{d-k}(T).$ (3.40)
This implies that $f\in\raisebox{1.0pt}{$\scriptstyle\bigstar$}W^{d-k}(T)$. But in view of (3.8) we have $\raisebox{1.0pt}{$\scriptstyle\bigstar$}W^{d-k}(T)\cap\mathcal{P}^{r-1}\Lambda^{k}(T)=0$, so we get $f=0$.
@@
Any element $u\in A^{k}(T)$ is uniquely determined by the following degrees of freedom, where $T^{\prime}$ are faces of $T$:
$\int_{T^{\prime}}u\wedge e,\quad e\in\mathcal{P}_{\!\!-}^{r}\Lambda^{d^{\prime}-k}(T^{\prime}),\ d^{\prime}=\dim T^{\prime}.$ (3.41)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=\int_{T}f\cdot u^{\prime},$ (3.43)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=\int_{T}g\cdot v^{\prime}.$ (3.44)
@@
$\int_{T}u\wedge e=0.$ (3.45)
Using (3.8) we then get, actually, that for all $e\in\mathcal{P}_{\!\!-}^{r+1}\Lambda^{d-k}(T)$:
$\int_{T}u\wedge e=0.$ (3.46)
@@
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=0,$ (3.48)
$\int_{T}g\cdot v^{\prime}$ $=\int_{T}g\cdot\mathrm{d}v^{\prime\prime}=0.$ (3.49)
Combining this with (3.44) we conclude that for all $v^{\prime}\in Z^{k-1}_{0}(T)$:
$\int_{T}u\cdot\mathrm{d}v^{\prime}=\int_{T}g\cdot v^{\prime}.$ (3.50)
@@
$\tilde{Z}^{k}_{0}(T)=\{u\in X^{k}_{0}(T)\ :\ \forall e\in\mathcal{P}_{\!\!-}^{r+1}\Lambda^{d-k}(T)\quad\int_{T}u\wedge e=0\}.$ (3.51)
We have $\tilde{Z}^{k}_{0}(T)\subseteq Z^{k}_{0}(T)$. The complex $\tilde{Z}^{\scriptscriptstyle{\bullet}}_{0}(T)$ is exact, by the snake lemma and the exactness properties of $\mathcal{P}_{\!\!-}^{r+1}\Lambda^{{\scriptscriptstyle{\bullet}}}(T)$. It is a variant of Proposition 2.1 but now exactness of $\tilde{Z}^{\scriptscriptstyle{\bullet}}_{0}(T)$ holds also at index $k=d$.
Since $\mathrm{d}u=0$ and (3.46) holds we deduce that there is $t\in\tilde{Z}^{k-1}_{0}(T)$ such that $\mathrm{d}t=u$. This entails that for all $e\in\mathcal{P}_{\!\!-}^{r+1}\Lambda^{d-k+1}(T)$:
$\int_{T}t\wedge e=0.$ (3.52)
Then we have, setting $u=\mathrm{d}t$ and $v^{\prime}=t$ in (3.50):
$\int_{T}\mathrm{d}t\cdot\mathrm{d}t=\int_{T}g\cdot t=\int_{T}t\wedge\raisebox{1.0pt}{$\scriptstyle\bigstar$}g=0.$ (3.53)
@@
We have proved that the bilinear form (3.42) defines an injective map:
$A^{k}_{0}(T)\to\mathcal{P}_{\!\!-}^{r}\Lambda^{d-k}(T)^{\star}.$ (3.54)
@@
For $k<d$, we have from [22, Equation (2.28)]:
$\dim G^{k}(T)=\dim\mathcal{P}_{\!\!-}^{r}\Lambda^{d-k}(T).$ (3.55)
@@
$\dim A^{d}(T)=\dim\mathcal{P}^{r}\Lambda^{d}(T)=\dim\mathcal{P}_{\!\!-}^{r}\Lambda^{0}(T).$ (3.57)
It follows that the bilinear form (3.42) is invertible.
(v) Now it follows from Proposition 3.5 that the FES admits extensions in the sense of [28, Definition 2.3] (or [27, Definition page 75]). The unisolvence of the degrees of freedom follows then from a general induction argument given for FES in [28, Proposition 2.5] (or [27, Proposition 5.35]). ∎
@@
For elements $u\in A^{k}(T)$ its moments against $\mathcal{P}^{r}\Lambda^{d-k}(T)$ can be computed inductively from the degrees of freedom.
@@
$(d-k+r)w=\mathrm{d}\kappa_{T}w+\kappa_{T}\mathrm{d}w-Q_{r-1}w.$ (3.59)
@@
$\int_{T}u\wedge\mathrm{d}\kappa_{T}w=\pm\int_{\partial T}u\wedge\kappa_{T}w\pm\int_{T}\mathrm{d}u\wedge\kappa_{T}w.$ (3.60)
@@
$\int_{T^{\prime}}u\wedge(\kappa_{T}-\kappa_{T^{\prime}})w+\int_{T^{\prime}}u\wedge(\kappa_{T^{\prime}}w-P_{r}\kappa_{T^{\prime}}w)+\int_{T^{\prime}}u\wedge P_{r}\kappa_{T^{\prime}}w.$ (3.61)
@@
In (3.61) the first and last terms are moments against $\mathcal{P}^{r}\Lambda^{d-1-k}(T^{\prime})$, so amenable to an induction step. The middle term is zero, since it is a moment against $U^{d-1}(T^{\prime})$ representing a constraint.
@@
(iii) In the general case we remark that $\mathrm{d}\mathrm{d}u=0$ and that the degrees of freedom of $\mathrm{d}u$ are computable from those of $u$. Applying the intermediate step to $\mathrm{d}u$, its moments against polynomials of degree $r$ can be computed. Going back to (3.60) and looking at the last term, we write:
$\int_{T}\mathrm{d}u\wedge\kappa_{T}w=\int_{T}\mathrm{d}u\wedge(\kappa_{T}w-P_{r}\kappa_{T}w)+\int_{T}\mathrm{d}u\wedge P_{r}\kappa_{T}w.$ (3.62)